Solve the initial value problem below using the method of Laplace transforms. y + 4y + 3y = 45 e 21, y(0) = -6, y(0) = 21
Answer #1

Page ① : Cuen, y + 4y+ 3y = 45 e2t, yco) = -6 , yCO) = 21 Applying Laplace bransforms on both sides, we get, L {y + 4y + 3Page ② : → (S+8) cs+2) L {y} = 45 - (65+3) 95 to S-2 45 (5+3) CS+2) CS-2) 6543 CS+3) CS+2) 3 L{y} = 45.- C6S+3) CS-2) (st3) C• Page ③ : = -30 = A (10) > A=-3 Puthing. 5=-, in eq @ we get, - 6 (1) + 9 (-1) + 51 = B (2) (-3) 3 36 = -6B1 B = -6 PuttingPage 6 :- Applying inverse Laplace on both sides, we get y - 4 s 6 1-2 = L 14 -3 1. (St3 2 J s 3 (S-2 S+1 By linearity, we ha